# UMA: Ultra-detailed Human Avatars via Multi-level Surface Alignment

### v1

@@ Abstract
Learning an animatable and clothed human avatar model with vivid dynamics and photorealistic appearance from multi-view videos is an important foundational research problem in computer graphics and vision. Fueled by recent advances in implicit representations, the quality of the animatable avatars has achieved an unprecedented level by attaching the implicit representation to drivable human template meshes. However, they usually fail to preserve highest level of detail, particularly apparent when the virtual camera is zoomed in and when rendering at 4K resolution and higher. We argue that this limitation stems from inaccurate surface tracking, specifically, depth misalignment and surface drift between character geometry and the ground truth surface, which forces the detailed appearance model to compensate for geometric errors. To address this, we propose a latent deformation model and supervising the 3D deformation of the animatable character using guidance from foundational 2D video point trackers, which offer improved robustness to shading and surface variations, and are less prone to local minima than differentiable rendering. To mitigate the drift over time and lack of 3D awareness of 2D point trackers, we introduce a cascaded training strategy that generates consistent 3D point tracks by anchoring point tracks to the rendered avatar, which ultimately supervise our avatar at vertex and texel level. To validate the effectiveness of our approach, we introduce a novel dataset comprising five multi-view video sequences, each over 10 minutes in duration, captured using 40 calibrated 6K-resolution cameras, featuring subjects dressed in clothing with challenging texture patterns and wrinkle deformations. Our approach demonstrates significantly improved performance in rendering quality and geometric accuracy over the prior state of the art.
Index Terms:
Human rendering, performance capture, 3D geometry and appearance synthesis.
1 Introduction
Creating photorealistic, animatable full-body humans – i.e. a model that maps skeletal motion to deforming geometry and surface appearance – is a longstanding and challenging problem in computer graphics and vision. However, digitizing high-quality virtual doubles while preserving clothing dynamics and rendering them in novel poses typically requires substantial manual efforts from skilled artists. Therefore, automating the digitization of photorealistic virtual humans by learning from multi-view videos have consistently emerged as highly active research topics. The ultimate goal is to replicate the finest appearance details and dynamic surface deformations just like those in real-captured videos.
@@ 1 Introduction
Figure 1: Given skeletal poses and a virtual camera, UMA renders ultra-detailed clothed human appearance and synthesizes high-fidelity geometry. Additionally, we introduce a new dataset featuring multi-view 6K video recordings, capturing subjects wearing clothing with challenging texture patterns and rich dynamics.
@@
Driven by recent advances in neural 3D representations  [1, 2], the fidelity of animatable clothed human avatars [3, 4] has significantly improved. They typically attach neural representations onto a drivable human skeleton [5], a parametric human body model [6], or a deformable template mesh [7]. However, these methods still fall short in capturing and synthesizing high-frequency details, e.g., fine texture patterns. We argue that it is primarily due to imprecise surface tracking during training as many approaches either use body models [6] (not accounting for the surface deformations at all) or deformable templates whose deformations are not correctly tracked [4]. This work attempts to solve the inaccuracy in the template tracking – more specifically, the misalignment in depth and drifting on the surface – by jointly considering multi-level surface registration and neural human rendering, resulting in high-quality geometry as well as intricate appearance details.
@@
We found the inaccurate surface tracking partly arises due to fact that the skeletal motion alone is an insufficient input conditioning to learn surface deformations, i.e., clothing dynamics, since these are inherently stochastic and subject to other factors such as rest state. Interestingly, this stochasticity can be mostly observed and learned from longer training videos, which prior work often avoids or their results show a significant quality degradation when the number of training frames is increased. To mitigate this issue, we introduce a per-frame latent code to the geometry network accounting for the inherent one-to-many mapping ambiguity [8] between the surface dynamics and skeletal motion.
@@
Moreover, prior works often supervise the learned animatable template on 3D pointclouds [9], which suffers from drifting on the surface across frames. To account for this, some works [10] leverage differentiable rendering for supervision, which is known to get stuck in local minima. Instead, we propose to leverage a foundational point tracker [11], which computes correspondences among consecutive video frames. However, directly tracking correspondence over long videos remains challenging due to the inevitable accumulation of drift over time. To this end, we propose an avatar-guided point tracking strategy, which estimates correspondence between the template mesh and multi-view videos, and then supervises the surface dynamics. Moreover, we propose a visibility-aware, cascaded filtering strategy based on the animatable template to aggregate multi-view 2D correspondences and refine them into reliable 3D correspondences, which serve as a more precise supervision for the animatable clothed human geometry at each training stage.
@@
We found that most existing multi-view datasets of clothed humans are limited by image resolution, lack of challenging texture patterns, and restricted skeletal pose diversity. Towards better benchmarking, we present a novel dataset featuring five clothed human subjects, each recorded for ten minutes at 6K resolution using 40 calibrated cameras greatly sampling skeletal pose and induced surface dynamics. Our contributions are summarized as follows:
@@
Our qualitative and quantitative results (see Fig. 1) demonstrates a clear improvement over prior work in terms of detail recovery, dynamics capture, and overall visual appearance. For more video results, we also refer to the video link.
2 Related Work
@@
Our work focuses on animatable human rendering and geometry generation, where the model takes solely skeletal motions as input at test time. We do not cover methods related to replay [12, 13, 14, 15, 16, 17], reconstruction [18, 19, 20, 21, 22, 23, 24, 25, 26], or image-based free-view rendering [27, 28, 29, 9, 30]. In the following sections, we review related works on animatable avatars, categorized by their shape representations, namely, mesh-based, implicit-based, and point-based approaches.
@@
Mesh-based Approaches. Textured meshes are the most prevalent representation for modeling clothed human avatars due to their compatibility with existing rendering and animation pipelines. Early approaches reconstruct and animate the person-specific textured template through physical simulation [31, 32], retrieval from a video database [33], or interpolating within a texture stack [34, 35]. More recently, some works [36, 18, 37] adopt neural networks to learn the motion-dependent texture of the human template mesh from the multi-view videos. Habermann et al. [10] further models motion-dependent surface deformations using a learnable embedded graph  [38] and captures dynamic appearance through motion-aware texture maps. MeshAvatar [39] estimates motion-dependent surface deformations using front-and-back feature maps and jointly learns material properties from multi-view videos in an end-to-end fashion. However, bounded by the limited resolution of the template meshes, the mesh-based approaches usually fail to recover the fine-grained geometry and appearance details.
@@
Implicit-based Approaches. To increase the representation’s capacity for modeling detailed appearance and geometry, implicit-based approaches combine implicit fields with explicit shape proxies, i.e., virtual bones [5], parametric body models [40, 41, 42, 43], or person-specific template meshes [10, 21]. To better model the pose-dependent appearance of humans, recent studies  [8, 44, 45, 46, 47, 48, 4] further introduce motion-aware residual deformations in the canonical space on top of the mesh template. Neural Actor [8] leverages the texture space of a parametric human body model to extract pose-aware features for inferring motion-dependent dynamics, but it struggles to represent humans wearing loose clothing. Moreover, it usually requires seconds to render a frame, due to the extensive sampling needed and the large MLP used to parameterize the neural radiance field. TriHuman [49] represents the animatable character using a triplane defined over the texture space of the person-specific template mesh, enabling real-time photorealistic rendering and high-quality surface geometry generation at test time. However, prior approaches either overlook surface deformations that cannot be solely attributed to skeletal poses, or they suffer from surface drift, which evitable leads to blurred appearance details and over-smoothed wrinkle deformations.
@@
Point-based Approaches. Point clouds have long served as a powerful and widely adopted representation for human avatar modeling due to their high flexibility in terms of modeling geometric shapes. SCALE [50] and POP [51] learn non-rigid deformations of dynamic clothing by representing the clothed body as dense point clouds parameterized in the UV space of parametric body models [52]. To address the discontinuities in UV parameterization, FITE [53] extracts pose-dependent features from orthographic projective maps. In recent years, 3D Gaussian Splatting [2] has attracted considerable attention for its capability to generate high-quality renderings in real time, making it a popular choice for animatable clothed human avatars. GART [54], 3DGS-Avatar [55], GauHuman [56], and HUGS [57] learn a canonicalized clothed human body represented with 3D Gaussian Splats from monocular videos and animate it using linear blend skinning. However, since the attributes of the 3D Gaussian Splats are predicted by MLPs, they struggle to capture high-frequency appearance details due to the inherent smoothness bias of MLPs. Therefore, ASH [7], GaussianAvatar [58], and Animatable Gaussians [59] learn animatable characters with motion-aware appearance by leveraging convolutional neural networks in the UV space or in the orthographic projection space. While achieving notable improvements in rendering quality, they still fall short in capturing the finest level of appearance details and reconstructing coherent, detailed surfaces, due to surface drift. By establishing accurate correspondences between the generated clothed human surface and the multi-view video using 2D point tracks, our approach captures fine-grained appearance details and reconstructs detailed as well as space-time coherent surface dynamics.
3 Methodology
Figure 2: Overview. Our method, UMA, takes skeletal motion and the camera view as input and generates high-fidelity geometry and appearance. Based on the observation that inaccurate surface tracking would significantly hinder the learning of detailed appearances of the animatable avatars, we propose a latent deformation model an supervising the 3D deformation of the animatable character using guidance from foundational 2D video point trackers.
@@ 3 Methodology
We aim to learn a photorealistic clothed human avatar with ultra-detailed appearance and surface geometry from multi-view videos. Specifically, we target at capturing skeletal motion-aware surface dynamics and preserving finest details such as clothing wrinkles and texture pattens, allowing users to zoom in closely without noticeable quality degradation. To this end, we introduce UMA , Ultra-detailed animatable clothed human avatar with Multi-level surface Alignment, which during inference takes skeletal motion and camera views as input and generates photorealistic renderings at a resolution of $1620\times 3072$ and detailed geometry at $11$ fps. An overview of our approach is shown in Fig. 2.
@@
In the following section, we will first introduce the underlying representation for the photoreal avatar (Sec. 3.1) and the key technical challenge (Sec. 3.2). Subsequently, we elaborate how UMA captures fine-grained appearance and geometry details by improving surface tracking in terms of alignment in depth (Sec. 3.3) as well as surface drift at triangle (Sec. 3.4) and texel resolution (Sec. 3.5).
3.1 Gaussian-based Avatar Representation
@@
Data Assumptions. We assume a set of segmented multi-view videos $\mathbf{I}_{f,c}\in\mathbb{R}^{H\times W}$, where $f$ and $c$ denote the frame and camera indices, respectively. $W$ and $H$ correspond to the width and height of captured frames. Each frame $\mathbf{I}_{f,c}$ is annotated with the 3D skeletal pose $\bm{\theta}_{f}\in\mathbb{R}^{D}$ using a commercial markerless motion capture system [60], where $D$ denotes the degrees of freedom (DoFs) of the 3D skeletal pose. The motion $\bm{\bar{\theta}}_{f}\in\mathbb{R}^{k\times D}$ is derived from skeletal poses of a sliding window ranging from frame $f-k+1$ to frame $f$ where the root joint translation is normalized w.r.t. the $f$th frame. Moreover, we adopt implicit surface reconstruction [61] to recover per-frame ground truth geometry $\mathbf{V}_{f}^{\mathrm{gt}}$ and render it into depth maps $\mathbf{N}_{f,c}\in\mathbb{R}^{H\times W}$ for all camera views. Notably, the recovered ground truth surfaces are per-frame reconstructions and they lack temporal surface correspondence.
Drivable Human Template Mesh. We define a drivable and deformable template mesh
$M(\bm{\bar{\theta}}_{f})$ $=f_{\mathrm{dq}}(f_{\mathrm{eg}}(\mathcal{F_{\mathrm{eg}}}(\bar{\bm{\theta}}_{f}),\bar{\mathbf{V}})+\mathcal{F_{\mathrm{delta}}}(\bar{\bm{\theta}}_{f}),\bm{\theta}_{f},\mathbf{S})$ (1)
$=f_{\mathrm{dq}}(f_{\mathrm{eg}}(\mathbf{A}_{f},\mathbf{T}_{f},\bar{\mathbf{V}})+\bm{\delta}_{f},\bm{\theta}_{f},\mathbf{S})$ (2)
$=\mathbf{V}_{f}$ (3)
@@
of a clothed human to model coarse-level geometry. It takes the skeletal motion $\bar{\bm{\theta}}_{f}$ as input and regresses posed and non-rigidly deformed 3D vertices $\mathbf{V}_{f}$ of a person-specific template mesh $\bar{\mathbf{V}}\in\mathbb{R}^{V\times 3}$. Precisely, to model the motion-aware deformation of the clothed human, we follow Habermann et al.[10] and first apply embedded deformation $f_{\mathrm{eg}}(\cdot)$ [38] where deformation parameters are predicted by a graph convolutional neural network $\mathcal{F_{\mathrm{eg}}}(\bar{\bm{\theta}}_{f})=\mathbf{A}_{f},\mathbf{T}_{f}$. Further, per-vertex displacements are predicted again as a function of skeletal motion by a second network $\mathcal{F_{\mathrm{delta}}}(\bar{\bm{\theta}}_{f})=\bm{\delta}_{f}$. Those deformations are applied onto the pose canonical mesh template $\bar{\mathbf{V}}$ from coarse to fine, i.e., first embedded deformations are applied followed by the displacements. Lastly, the canonical and non-rigidly deformed template $\bar{\mathbf{V}}_{f}$ is posed w.r.t. the skeletal pose ${\bm{\theta}}_{f}$ using Dual Quaternion skinning [62] $f_{\mathrm{dq}}(\cdot)$ defining $\mathbf{V}_{f}$. Here, $\mathbf{S}$ denotes the skinning weights matrix.
@@
$\mathcal{L}_{\mathrm{template}}(\mathbf{V}_{f})=\mathcal{L}_{\mathrm{cham}}(\mathbf{V}_{f},\mathbf{V}^{\mathrm{gt}}_{f})+\mathcal{L}_{\mathrm{spatial}}(\mathbf{V}_{f})$ (4)
@@
for all frames, where the first term compares the posed and deformed template against the ground truth surface in terms of Chamfer distance while the second term is a combination of spatial regularizers.

### v2

@@ Abstract.
Learning an animatable and clothed human avatar model with vivid dynamics and photorealistic appearance from multi-view videos is an important foundational research problem in computer graphics and vision. Fueled by recent advances in implicit representations, the quality of the animatable avatars has achieved an unprecedented level by attaching the implicit representation to drivable human template meshes. However, they usually fail to preserve highest level of detail, e.g., fine textures and yarn-level patterns, particularly apparent when the virtual camera is zoomed in and when rendering at 4K resolution and higher. We argue that this limitation stems from inaccurate surface tracking, specifically, depth misalignment and surface drift between character geometry and the ground truth surface, which forces the detailed appearance model to compensate for geometric errors. To address this, we adopt a latent deformation model and supervise the 3D deformation of the animatable character using guidance from foundational 2D video point trackers, which offer improved robustness to shading and surface variations, and are less prone to local minima than differentiable rendering. To mitigate the drift over time and lack of 3D awareness of 2D point trackers, we introduce a cascaded training strategy that generates consistent 3D point tracks by anchoring point tracks to the rendered avatar, which ultimately supervise our avatar at vertex and texel level. Furthermore, a lightweight Gaussian texture super-resolution module is employed to reconstruct challenging appearance details and micro-level structures using localized information. To validate the effectiveness of our approach, we introduce a novel dataset comprising five multi-view video sequences, each over 10 minutes in duration, captured using 40 calibrated 6K-resolution cameras, featuring subjects dressed in clothing with challenging texture patterns and wrinkle deformations. Our approach demonstrates significantly improved performance in rendering quality and geometric accuracy over the prior state of the art.
Human rendering, performance capture.
$c$s: Computing methodologies Renderingccs: Computing methodologies Computer visionccs: Computer systems organization Real-time systems
@@
Figure 1. Given skeletal poses and a virtual camera, UMA renders ultra-detailed clothed human appearance and synthesizes high-fidelity geometry. Notably, UMA enables users to digitally zoom in, allowing close inspection of texture details or even fine yarn-level patterns. Additionally, we introduce a new dataset featuring multi-view 6K video recordings, capturing subjects wearing clothing with challenging texture patterns and rich dynamics. The reconstructed animatable avatars can serve as a strong foundation for downstream applications, e.g., motion editing, pose re-targeting, and texture editing. The fidelity of the reconstructed avatars makes them particularly suitable for virtual and mixed reality, where users can closely observe fine-grained appearance details.
1. Introduction
Creating photorealistic, animatable full-body humans – i.e. a model that maps skeletal motion to deforming geometry and surface appearance – is a longstanding and challenging problem in computer graphics and vision. However, digitizing high-quality virtual doubles while preserving clothing dynamics and rendering them in novel poses typically requires substantial manual efforts from skilled artists (Wang et al., 2024). Therefore, automating the digitization of photorealistic virtual humans by learning from multi-view videos have consistently emerged as highly active research topics. The ultimate goal is to replicate the finest appearance details and dynamic surface deformations just like those in real-captured videos.
@@ 1. Introduction
Driven by recent advances in neural 3D representations  (Su et al., 2021; Kerbl et al., 2023), the fidelity of animatable clothed human avatars (Li et al., 2024; Habermann et al., 2023) has significantly improved. They typically attach neural representations onto a drivable human skeleton (Li et al., 2022), a parametric human body model (Wang et al., 2022), or a deformable template mesh (Pang et al., 2024). However, these methods still fall short in capturing and synthesizing high-frequency details, e.g., fine texture patterns and detailed geometry. We argue that it is primarily due to imprecise surface tracking during training as many approaches either use body models (Wang et al., 2022) (not accounting for the surface deformations at all) or deformable templates whose deformations are not correctly tracked (Habermann et al., 2023, 2021a), which leads to conflicting supervision across views and frames. This work attempts to solve the inaccuracy in the template tracking – more specifically, the misalignment in depth and drifting on the surface – by jointly considering multi-level surface registration and neural human rendering, resulting in noticeable improvements in recovering high-quality geometry and intricate appearance details.
@@
The misalignment in depth denotes the discrepancies between the learned and the ground-truth geometry along the camera view direction. We found the depth misalignment partly arises due to fact that the skeletal motion alone is an insufficient input conditioning to learn surface deformations, i.e., clothing dynamics, since these are inherently stochastic and subject to other factors such as rest state. Interestingly, this stochasticity can be mostly observed and learned from longer training videos, which prior work (Wang et al., 2022; Li et al., 2022) often avoids or their results show a significant quality degradation when the number of training frames is increased. To mitigate this issue, we adopt a per-frame latent code to a coarse geometry network accounting for the inherent one-to-many mapping ambiguity (Liu et al., 2021) between the surface dynamics and skeletal motion during training. This effectively builds a subspace for large-scale clothing dynamics, which better captures the cloth deformations present in the real video, and facilitates the learning of surface correspondence.
@@
Moreover, prior works often supervise the learned animatable template on 3D point-clouds (Shetty et al., 2024; Habermann et al., 2021a, 2023), which suffers from drifting on the surface across frames, due to the missing surface correspondence. To establish correspondence between the learned geometry and the ground truth surface observed in the multi-view video, some works (Habermann et al., 2021a) leverage differentiable rendering for supervision, which is known to get stuck in local minima due to the shading variations and the complicate clothing dynamics. Instead, we leverage a foundational point tracker (Karaev et al., 2024), which is originally designed to compute correspondences across video frames. However, directly tracking correspondence over long videos remains challenging due to the inevitable accumulation of drift over time. To this end, we propose an avatar-guided point tracking strategy, which estimates correspondence between the template mesh and multi-view videos, effectively supervising the 3D surface dynamics. Moreover, we propose a visibility-aware, cascaded filtering strategy based on the animatable template to aggregate multi-view 2D correspondences and refine them into reliable 3D correspondences, which serve as a more precise supervision for the animatable clothed human geometry at each training stage.
Although our method effectively resolves depth misalignment and surface drifting, achieving precise alignment between the dynamic surface of the animatable character and the ground-truth surface at the depth, vertex, and texel levels. However, it may still struggle to recover the most intricate appearance details, such as fine yarn patterns, due to the limited number of 3D Gaussian splats. To this end, we further introduce a lightweight super-resolution module that up-samples the dynamic Gaussian splats with minimal computational overhead.
@@
We found that most existing multi-view datasets of clothed humans are limited by image resolution  (Habermann et al., 2021b; Peng et al., 2021b), lack of challenging texture patterns  (Li et al., 2024), and restricted skeletal pose diversity  (Peng et al., 2021b). Towards better benchmarking, we present a novel dataset featuring five clothed human subjects, each recorded for ten minutes at 6K resolution using 40 calibrated cameras greatly sampling skeletal pose and induced surface dynamics. Our contributions are summarized as follows:
@@
Our qualitative and quantitative results (see Fig. 1) demonstrate a clear improvement over prior work in terms of detail recovery, dynamics capture, and overall visual appearance.
2. Related Work
@@
Our work focuses on animatable human rendering and geometry generation, where the model takes solely skeletal motions as input at test time. We do not cover methods related to replay (Peng et al., 2021b; Lombardi et al., 2021; Wang et al., 2020; Weng et al., 2022; Işık et al., 2023; Xu et al., 2024; Jiang et al., 2025), reconstruction (Xiang et al., 2021; Alldieck et al., 2018, 2019; Habermann et al., 2020, 2019; Xiu et al., 2023; Zhang et al., 2024; Zheng et al., 2025; Zhu et al., 2022), or image-based free-view rendering (Kwon et al., 2021; Wang et al., 2021b; Remelli et al., 2022; Shetty et al., 2024; Sun et al., 2025). In the following sections, we review related works on animatable avatars, categorized by their shape representations, namely, mesh-based, implicit-based, and point-based approaches.
@@
Mesh-based Approaches. Textured meshes are the most prevalent representation for modeling clothed human avatars due to their compatibility with existing rendering and animation pipelines. Early approaches reconstruct and animate the person-specific textured template through physical simulation (Stoll et al., 2010; Guan et al., 2012), retrieval from a video database (Xu et al., 2011), or interpolating within a texture stack (Casas et al., 2014; Shysheya et al., 2019). More recently, some works (Bagautdinov et al., 2021; Xiang et al., 2021, 2022) adopt neural networks to learn the motion-dependent texture of the human template mesh from the multi-view videos. Habermann et al. (2021a) further models motion-dependent surface deformations using a learnable embedded graph  (Sumner et al., 2007) and captures dynamic appearance through motion-aware texture maps. MeshAvatar (Chen et al., 2024) estimates motion-dependent surface deformations using front-and-back feature maps and jointly learns material properties from multi-view videos in an end-to-end fashion. However, bounded by the limited resolution of the template meshes, the mesh-based approaches usually fail to recover the fine-grained geometry and appearance details.
@@
Implicit-based Approaches. To increase the representation’s capacity for modeling detailed appearance and geometry, implicit-based approaches combine implicit fields with explicit shape proxies, i.e., virtual bones (Li et al., 2022), parametric body models (Loper et al., 2015a; Osman et al., 2020; Pavlakos et al., 2019; Joo et al., 2018), or person-specific template meshes (Habermann et al., 2021a, 2020). To better model the pose-dependent appearance of humans, recent studies  (Liu et al., 2021; Peng et al., 2021a; Xu et al., 2021; Gao et al., 2023; Zheng et al., 2023; Kwon et al., 2023; Habermann et al., 2023) further introduce motion-aware residual deformations in the canonical space on top of the mesh template. Neural Actor (Liu et al., 2021) leverages the texture space of a parametric human body model to extract pose-aware features for inferring motion-dependent dynamics, but it struggles to represent humans wearing loose clothing. Moreover, it usually requires seconds to render a frame, due to the extensive sampling needed and the large MLP used to parameterize the neural radiance field. TriHuman (Zhu et al., 2023) represents the animatable character using a triplane defined over the texture space of the person-specific template mesh, enabling real-time photorealistic rendering and high-quality surface geometry generation at test time. However, prior approaches either overlook surface deformations that cannot be solely attributed to skeletal poses, or they suffer from surface drift, which evitable leads to blurred appearance details and over-smoothed geometry.
@@
Point-based Approaches. Point clouds have long served as a powerful and widely adopted representation for human avatar modeling due to their high flexibility in terms of modeling geometric shapes. SCALE (Ma et al., 2021a) and POP (Ma et al., 2021b) learn non-rigid deformations of dynamic clothing by representing the clothed body as dense point clouds parameterized in the UV space of parametric body models (Loper et al., 2015b). To address the discontinuities in UV parameterization, FITE (Lin et al., 2022) extracts pose-dependent features from orthographic projective maps. In recent years, 3D Gaussian Splatting (Kerbl et al., 2023) has attracted considerable attention for its capability to generate high-quality renderings in real time, making it a popular choice for animatable clothed human avatars. GART (Lei et al., 2024), 3DGS-Avatar (Qian et al., 2024), GauHuman (Hu and Liu, 2024), and HUGS (Kocabas et al., 2024) learn a canonicalized clothed human body represented with 3D Gaussian Splats from monocular videos and animate it using linear blend skinning. However, since the attributes of the 3D Gaussian Splats are predicted by MLPs, they struggle to capture high-frequency appearance details due to the inherent smoothness bias of MLPs. Therefore, ASH (Pang et al., 2024), GaussianAvatar (Hu et al., 2024), and Animatable Gaussians (Li et al., 2024) learn animatable characters with motion-aware appearance by leveraging convolutional neural networks in the UV space or in the orthographic projection space. While achieving notable improvements in rendering quality, they still fall short in capturing the finest level of appearance details and reconstructing coherent, detailed surfaces, due to surface drift. PhysAvatar (Zheng et al., 2024) tracks the human and clothing surface by initializing 3DGS on the template mesh and supervising the deformation with a photometric loss. However, since the color of the Gaussians is optimizable over time, the geometric error can be compensated with color changes, resulting in wrong correspondences and, therefore, blurred appearance details. Moreover, generating the simulation ready clothing requires immense efforts from experts to resolve the collisions, and drape the garments on the body template at the initialization phase. In contrast, by establishing accurate correspondences between the generated clothed human surface and the multi-view video using 2D point tracks, our approach captures fine-grained appearance details and reconstructs detailed as well as space-time coherent surface dynamics without manual intervention from the artists.
3. Methodology
Figure 2. Overview. UMA, takes skeletal motion and the camera view as input and generates high-fidelity geometry and appearance. To enhance the fidelity of the reconstructed human appearance and geometry, we tackle the problem from two key perspectives: avatar representation and multi-level surface alignment. For avatar representation, to address the stochasticity of the clothing dynamics that cannot be modeled by the skeletal motions, we inject a learnable latent code $\mathbf{z}_{f}$(zero latent $\mathbf{z}_{0}$ for testing) the drivable template $\mathbf{V}_{f}$ (Sec. 3.3). A texel super resolution module $\mathcal{E}_{\mathrm{sr}}$ is adopted to densify the animatable gaussian textures(Sec. 3.6). For multi-level surface alignment, we supervise the surface geometry at both the vertex (Sec. 3.4) and texel levels (Sec. 3.5) using novel supervision derived from a foundational 2D point tracker. Specifically, the 2D point tracks $\mathbf{P}_{f,c,i}$ between the rasterized and ground-truth images obtained from the tracker are lifted and aggregated into 3D correspondences $\tilde{\mathbf{P}}_{f,i}$ across multiple views using the drivable template $\mathbf{V}_{f}$.
@@ 3. Methodology
We aim to learn a photorealistic clothed human avatar with ultra-detailed appearance and surface geometry from multi-view videos. Specifically, we target at capturing skeletal motion-aware surface dynamics and preserving finest details such as clothing wrinkles and texture pattens, allowing users to zoom in closely without noticeable quality degradation. To this end, we introduce UMA , Ultra-detailed animatable clothed human avatar with Multi-level surface Alignment, which during inference takes skeletal motion and camera views as input and generates photorealistic renderings at a resolution of $1620\times 3072$ and detailed geometry $18$ fps. An overview of our approach is illustrated in Fig. 2.
@@
In the following section, we will first introduce the underlying representation for the photoreal avatar (Sec. 3.1) and the key technical challenge (Sec. 3.2). Subsequently, we elaborate how UMA captures fine-grained appearance and geometry details by improving surface tracking in terms of alignment in depth (Sec. 3.3) as well as surface drift at vertex (Sec. 3.4) and texel level (Sec. 3.5). Finally, we present the light-weight texel super-resolution module for modeling finest appearance details (Sec. 3.6). The adopted supervision and trained modules for each stage are summarized in Tab. 1.
3.1. Gaussian-based Avatar Representation
@@
Data Assumptions. We assume segmented multi-view videos $\mathbf{I}_{f,c}\in\mathbb{R}^{H\times W}$, where $f$ and $c$ denote the frame and camera indices, respectively. $W$ and $H$ denotes the width and height of captured multi-view imagery. Each frame $\mathbf{I}_{f,c}$ is annotated with the camera camera calibrations $\mathbf{C}_{c}$ and 3D skeletal pose $\boldsymbol{\theta}_{f}\in\mathbb{R}^{D}$ using a commercial markerless motion capture system (TheCaptury, 2020), where $D$ denotes the degrees of freedom (DoFs) of the 3D skeletal pose. The motion $\boldsymbol{\bar{\theta}}_{f}\in\mathbb{R}^{k\times D}$ is derived from skeletal poses of a sliding window ranging from frame fk+1f-k+1 to frame $f$ where the root joint translation is normalized w.r.t. the $f$th frame. Moreover, we adopt implicit surface reconstruction (Wang et al., 2023) to recover per-frame ground truth geometry $\mathbf{V}_{f}^{\mathrm{gt}}$ and render it into depth maps $\mathbf{N}_{f,c}\in\mathbb{R}^{H\times W}$ for all camera views. Notably, the reconstructed ground truth surfaces $\mathbf{V}_{f}^{\mathrm{gt}}$ are per-frame reconstructions and therefore they lack surface correspondence over time.
Drivable Human Template Mesh. We first define a drivable and deformable template mesh
(1) $M(\boldsymbol{\bar{\theta}}_{f})$ $=f_{\mathrm{dq}}(f_{\mathrm{eg}}(\mathcal{F_{\mathrm{eg}}}(\bar{\boldsymbol{\theta}}_{f}),\bar{\mathbf{V}})+\mathcal{F_{\mathrm{delta}}}(\bar{\boldsymbol{\theta}}_{f}),\boldsymbol{\theta}_{f})$
(2) $=f_{\mathrm{dq}}(f_{\mathrm{eg}}(\mathbf{A}_{f},\mathbf{T}_{f},\bar{\mathbf{V}})+\boldsymbol{\delta}_{f},\boldsymbol{\theta}_{f})$
(3) $=f_{\mathrm{dq}}(\bar{\mathbf{V}}_{f},\boldsymbol{\theta}_{f})$
(4) $=\mathbf{V}_{f}$
@@
of a clothed human to model coarse-level geometry. It takes the skeletal motion $\bar{\boldsymbol{\theta}}_{f}$ as input and regresses posed and non-rigidly deformed 3D vertices $\mathbf{V}_{f}$ of a person-specific template mesh $\bar{\mathbf{V}}\in\mathbb{R}^{V\times 3}$. Precisely, to model the motion-aware deformation of the clothed human, we follow  Habermann et al. (2021a) and first apply embedded deformation $f_{\mathrm{eg}}(\cdot)$ (Sumner et al., 2007) where deformation parameters are predicted by a graph convolutional neural network
(5) $\mathcal{F_{\mathrm{eg}}}(\bar{\boldsymbol{\theta}}_{f})=\mathbf{A}_{f},\mathbf{T}_{f},$
where $\mathbf{A}_{f}\in\mathbb{R}^{V\times 4}$ and $\mathbf{T}_{f}\in\mathbb{R}^{V\times 4}$ denotes the translation and rotation quaternions for the embedded graph nodes in the canonical space. Further, per-vertex displacements $\boldsymbol{\delta}_{f}$ are predicted as a function of skeletal motion by a second network
(6) $\boldsymbol{\delta}_{f}=\mathcal{F_{\mathrm{delta}}}(\bar{\boldsymbol{\theta}}_{f}).$
Those deformations are applied onto the canonical mesh template $\bar{\mathbf{V}}$ from coarse to fine, i.e., first embedded deformations are applied followed by the per-vertex displacements. Lastly, the canonical and non-rigidly deformed template $\bar{\mathbf{V}}_{f}$ is posed w.r.t. the skeletal pose ${\boldsymbol{\theta}}_{f}$ using Dual Quaternion skinning (Kavan et al., 2007) $f_{\mathrm{dq}}(\cdot)$.
@@
(7) $\mathcal{L}_{\mathrm{temp}}(\mathbf{V}_{f})=\mathcal{L}_{\mathrm{cham}}(\mathbf{V}_{f},\mathbf{V}^{\mathrm{gt}}_{f})+\mathcal{L}_{\mathrm{spatial}}(\mathbf{V}_{f})$
@@
for all frames, where the first term compares the posed and deformed template against the ground truth surface in terms of Chamfer distance while the second term
(8) $\mathcal{L}_{\mathrm{spatial}}(\mathbf{V}_{f})=\mathcal{L}_{\mathrm{lap}}(\mathbf{V}_{f})+\mathcal{L}_{\mathrm{lapz}}(\mathbf{V}_{f})+\mathcal{L}_{\mathrm{norm}}(\mathbf{V}_{f})$
is a combination of spatial regularization terms, namely, the mesh Laplacian loss $\mathcal{L}_{\mathrm{lap}}$, Laplacian smoothness term $\mathcal{L}_{\mathrm{lapz}}$, and face normal consistency loss $\mathcal{L}_{\mathrm{norm}}$. We refer to the supplemental document for more details regarding the embedded deformation and the geometry regularization.